OpenStudy (anonymous):

Factor completely: 36x2 − 49 (6x − 1)(6x − 49) (6x + 1)(6x − 49) (6x + 7)(6x − 7) (6x − 7)(6x − 7)

OpenStudy (anonymous):

\(a^2-b^2=(a+b)(a-b) \) in your case \(a=6x,b=7\)

OpenStudy (anonymous):

I kept getting that it cannot be factored...

OpenStudy (anonymous):

So it would be (6x + 7)(6x − 7)?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

(6x)^2 - 7^2 think of it like that as @satellite73 said

OpenStudy (anonymous):

your job was to recognize that \(36x^2=(6x)^2\) and that \(49=7^2\)